# Kinetic Theory of Quasiparticles, Retarded Correlators and Hydrodynamics

### v1

@@ III Charge Diffusion Correlators
$G^{(s)}_{00}(x;\omega,\vec{k})=\chi(x)\frac{1+\frac{i\omega\tau_{eq}-1}{2i\tau_{eq}k\chi(x)}\mathcal{S}^{(s)}(x;\omega,\vec{k})}{1-\frac{\mathcal{S}^{(s)}(x;\omega,\vec{k})}{2ik\tau_{eq}\chi(x)}},$ (27a)
$G^{(s)}_{03}(x;\omega,\vec{k})=\frac{\omega}{k}G^{(s)}_{00}(x;\omega,\vec{k}),$ (27b)
$G^{(s)}_{33}(x;\omega,\vec{k})=\frac{\omega^{2}}{k^{2}}G^{(s)}_{00}(x;\omega,\vec{k}),$ (27c)
$G^{(s)}_{11}(x;\omega,\vec{k})=G^{(s)}_{22}(x;\omega,\vec{k})=-\frac{i\tau_{eq}\omega\chi(x)}{2}\left(\frac{1-i\tau_{eq}\omega}{(\tau_{eq}\,k)^{2}}+\frac{(1-i\tau_{eq}\omega)^{2}\mathcal{S}^{(s)}(x;\omega,\vec{k})}{2(i\tau_{eq}\,k)^{3}\chi(x)}+\frac{i\mathcal{T}^{(s)}(x;\omega,\vec{k})}{2k\tau_{eq}\chi(x)}\right).$ (27d)
where $s=(l,h)$. For a massless system of particles, i.e. $x\to 0$, we obtain the previous results [27].
@@ III.1 Hydrodynamic limit
$G^{(l)}_{00}(x;\omega,\vec{k})=\chi(x)\frac{\mathcal{N}^{0(l)}(x;\omega,\vec{k})}{\mathcal{D}^{0(l)}(x;\omega,\vec{k})},$ (28)
@@
$\mathcal{N}^{0(l)}(x;\omega,\vec{k})=\frac{i(1-i)(1-i\tau_{eq}\omega)}{\tau_{eq}\omega}-\frac{ik^{2}\tau_{eq}N(x)}{15\omega(1-i\tau_{eq}\omega)K_{2}(x)\left(x^{2}K_{0}(x)+xK_{1}(x)-K_{2}(x)\right)},$
$\mathcal{D}^{0(l)}(x;\omega,\vec{k})=1-\frac{ik^{2}\tau_{eq}D(x)}{15\omega(1-i\tau_{eq}\omega)^{2}K_{2}(x)\left(x^{2}K_{0}(x)+xK_{1}(x)-K_{2}(x)\right)},$ (29)
@@
Hydro modes are solutions of spectral curves $\mathcal{D}^{(l)}(x;\omega,\vec{k})=0$ at small momenta which are obtained as follows
$\omega^{(nh)}_{(1,2)}$ $=-\frac{i}{\tau_{eq}}\pm\frac{k}{\sqrt{15}}\sqrt{\frac{5x^{2}\left(xK_{0}(x)+K_{1}(x)\right)^{2}+xK_{2}(x)\left(17xK_{0}(x)+8K_{1}(x)\right)-4K^{2}_{2}(x)}{K_{2}(x)\left(x^{2}K_{0}(x)+xK_{1}(x)-K_{2}(x)\right)}}$
$+\frac{ik^{2}\tau_{eq}}{6}\left(1-\frac{x^{2}K_{0}(x)+xK_{1}(x)}{K_{2}(x)}\right)+\mathcal{O}(k^{3}),$
$\omega^{(h)}_{(3)}$ $=-\frac{ik^{2}\tau_{eq}}{3}\left(1-\frac{x^{2}K_{0}(x)+xK_{1}(x)}{K_{2}(x)}\right)+\mathcal{O}(k^{4}),$ (31)
@@
$G^{(h)}_{00}(x;\omega,\vec{k})\bigg{|}_{x\to\infty}=0.$ (32)
This implies that for extremely massive quasiparticles, the hydrodynamic modes are not accounted for. In other words, collective modes only emerge for light particles. In Fig. 2 we plot the diffusive hydro modes for different cases. The left figure indicates the modes at $x=0.1$ for different $k\tau_{eq}$ values. The right figure shows the same results for $k\tau_{eq}=0.1$ for different $x$ values. It can be seen that the mass moves the poles into the unstable zone. This is because mass changes the diffusion constant, thereby influencing the location of hydro poles. We discuss it in the following lines.
To ensure the stability of hydro modes at low momenta, we need to check
$\mbox{Im}(\omega)\leq 0.$ (33)
This condition in the diffusive mode leads to
$1.2342\displaystyle 1-\frac{x^{2}K_{0}(x)+xK_{1}(x)}{K_{2}(x)}\geq 0,\quad\Rightarrow\quad x<$ 1.2342 . (34)
This means that small masses preserve the stability of hydro modes. Moreover, for $x>0.734$ the non-hydro modes become purely imaginary, which is derived from the positivity of $k$ coefficient in $\omega_{1,2}^{(nh)}$. Therefore, the $x<0.734$ zone gives rise to stable hydro-diffusive and non-hydro propagating modes, while the $1.23420.734<x<$ 1.2342 region yields the stable hydro-diffusive and non-hydro-diffusive modes.
@@
$\lim_{\omega\rightarrow 0}\lim_{k\rightarrow 0}\frac{\omega}{k^{2}}{\rm Im}\,G_{00}(\omega,{\bf k})$ $=\chi D\,$
$\lim_{\omega\rightarrow 0}\lim_{k\rightarrow 0}\frac{1}{k^{2}}{\rm Re}\,G_{00}(\omega,{\bf k})$ $=\chi D\tau_{D}.$ (35)
Using the expressions given in Eqs. (28) and (III.1) we readily obtain
$D(x)=\frac{\tau_{eq}}{3}\left(1-\frac{x^{2}K_{0}(x)+xK_{1}(x)}{K_{2}(x)}\right),$
$\tau_{D}=\tau_{eq}.$ (36)
The condition $x<1.2342$ that is already derived is important to have a positive diffusion constant.
@@ IV Energy-Momentum Correlators
$\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)=\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}=\frac{T_{0}^{4}x^{2}}{2\pi^{2}}\bigg{(}x^{2}K_{0}(x)+5xK_{1}(x)+12K_{2}(x)\bigg{)},$
@@
$\delta T^{0i}=-\frac{\tau_{eq}}{d}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{0i}(\omega,\vec{k})\frac{\mathcal{I}_{4}(v,\omega,\vec{k})}{i\tau_{eq}\omega}+\Gamma^{0}_{3i}(\omega,\vec{k})\mathcal{I}_{6}(v,\omega,\vec{k})\bigg{\}},\,\,i=1,2,$
$\delta T^{33}=\frac{\delta T^{00}c_{s}^{2}(1-i\tau_{eq}\omega)}{k^{2}\tau_{eq}^{2}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,\vec{k})+\frac{\delta T^{03}}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{5}(v,\omega,\vec{k})$
$\hskip 21.33955pt-\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{00}(\omega,\vec{k})\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}}\mathcal{I}_{1}(v,\omega,\vec{k})+2\,\Gamma^{0}_{03}(\omega,\vec{k})\mathcal{I}_{5}(v,\omega,\vec{k})+\Gamma^{0}_{33}(\omega,\vec{k})\mathcal{I}_{7}(v,\omega,\vec{k})$
$\hskip 142.26378pt+(\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\mathcal{I}_{8}(v,\omega,\vec{k})\bigg{\}},$
$\delta T^{ii}=\frac{\delta T^{00}c_{s}^{2}}{i\tau_{eq}\omega(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{4}(v,\omega,\vec{k})+\frac{\delta T^{03}}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,\vec{k})$
$\hskip 21.33955pt-\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{00}(\omega,\vec{k})\frac{\mathcal{I}_{4}(v,\omega,\vec{k})}{i\tau_{eq}\omega}+2\,\Gamma^{0}_{03}(\omega,\vec{k})\mathcal{I}_{6}(v,\omega,\vec{k})+\Gamma^{0}_{33}(\omega,\vec{k})\mathcal{I}_{8}(v,\omega,\vec{k})$
$\hskip 142.26378pt+(2\Gamma^{0}_{ii}(\omega,\vec{k})+\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\mathcal{I}_{9}(v,\omega,\vec{k})\bigg{\}},\,\,\,i=1,2,$
$\delta T^{12}=-\,\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\Gamma^{0}_{12}(\omega,\vec{k})\mathcal{I}_{9}(v,\omega,\vec{k}),$
$\delta T^{3i}=\frac{\delta T^{0i}}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,\vec{k})$
$\hskip 21.33955pt-\,\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{0i}(\omega,\vec{k})\mathcal{I}_{6}(v,\omega,\vec{k})+\Gamma^{0}_{3i}(\omega,\vec{k})\mathcal{I}_{8}(v,\omega,\vec{k})\bigg{\}},\,\,i=1,2.$
@@
$G^{00,00}(\omega,\vec{k})=-\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\bigg{(}1+\frac{i\,b\,c_{s}^{2}\,k\,\tau_{eq}}{ac-b^{2}c_{s}^{2}}\bigg{)},$
$G^{00,03}(\omega,\vec{k})=-(\varepsilon_{eq}+P_{eq})\,\frac{i\,\tau_{eq}\,\omega\,b}{ac-b^{2}c_{s}^{2}},$
$G^{00,33}(\omega,\vec{k})=\frac{\omega}{k}G^{00,03}(\omega,\vec{k}),$
$G^{00,11}(\omega,\vec{k})=G^{00,22}(\omega,\vec{k})$
@@
$G^{03,00}(\omega,\vec{k})=\frac{\omega}{k}\left(G^{00,00}(\omega,\vec{k})+\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\right)=-(\varepsilon_{eq}+P_{eq})\,\frac{\,i\,\tau_{eq}\,\omega\,b}{ac-b^{2}c_{s}^{2}}=G^{00,03}(\omega,\vec{k}),$
$G^{03,03}(\omega,\vec{k})=\frac{\omega}{k}G^{00,03}(\omega,\vec{k})-\left(\varepsilon_{eq}+P_{eq}\right),$
$G^{03,33}(\omega,\vec{k})=\frac{\omega}{k}G^{03,03}(\omega,\vec{k}),$
$G^{03,11}(\omega,\vec{k})=G^{03,22}(\omega,\vec{k})=\frac{\omega}{k}\left(G^{00,11}(\omega,\vec{k})-\left(\varepsilon_{eq}+P_{eq}\right)\right),$
$G^{0i,0i}(\omega,\vec{k})=\frac{k}{\omega}G^{0i,3i}(\omega,\vec{k})=\frac{-i\tau_{eq}\,k}{d}\,\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,\vec{k}),\qquad i=1,2,$
$G^{12,12}(\omega,\vec{k})=-i\tau_{eq}\omega\,\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{9}(v,\omega,\vec{k}),$
$G^{3i,0i}(\omega,\vec{k})=\frac{k}{\omega}G^{3i,3i}(\omega,\vec{k})=\frac{\omega}{k}G^{0i,0i}(\omega,\vec{k}).$ (47)
We check the Ward identity throughout the latter relations. Below, we give a detailed expansion of the $b$ parameter
$b\bigg{|}_{x\ll 1}$ $=\frac{1}{ik\tau_{eq}c_{s}^{2}}\bigg{\{}1+\frac{i\omega\tau_{eq}-1}{2ik\tau_{eq}}\bigg{(}L(\omega,\vec{k})+g_{1}(x)\left(L(\omega,\vec{k})-\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$\hskip 49.79231pt-g_{2}(x)\left(L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(k^{2}\tau_{eq}^{2}-(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\mathcal{O}(\frac{x^{6}}{\tilde{p}^{6}})\bigg{)}\bigg{\}},$
$b\bigg{|}_{x\gg 1}$ $=0,$
@@
$g_{1}(x)\equiv\frac{x^{2}\left(xK_{1}(x)+K_{2}(x)\right)}{2\left(x^{2}K_{0}(x)+5xK_{1}(x)+12K_{2}(x)\right)},$
$g_{2}(x)\equiv\frac{x^{4}\left(K_{2}(x)-3K_{0}(x)\right)}{16\left(x^{2}K_{0}(x)+5xK_{1}(x)+12K_{2}(x)\right)}.$ (49)
Since Eq. (IV) reveals that energy-momentum correlators are insensitive to large masses, we restrict our analysis to the small-mass limit in the subsequent part.

### v2

@@ III Charge Diffusion Correlators
$G^{(a)}_{00}(x;\omega,\vec{k})=\chi(x)\frac{1+\frac{i\omega\tau_{eq}-1}{2i\tau_{eq}k\chi(x)}\mathcal{S}^{(a)}(x;\omega,\vec{k})}{1-\frac{\mathcal{S}^{(a)}(x;\omega,\vec{k})}{2ik\tau_{eq}\chi(x)}},$ (27a)
$G^{(a)}_{03}(x;\omega,\vec{k})=\frac{\omega}{k}G^{(a)}_{00}(x;\omega,\vec{k}),$ (27b)
$G^{(a)}_{33}(x;\omega,\vec{k})=\frac{\omega^{2}}{k^{2}}G^{(a)}_{00}(x;\omega,\vec{k}),$ (27c)
$G^{(a)}_{11}(x;\omega,\vec{k})=G^{(a)}_{22}(x;\omega,\vec{k})=-\frac{i\tau_{eq}\omega\chi(x)}{2}\left(\frac{1-i\tau_{eq}\omega}{(\tau_{eq}\,k)^{2}}+\frac{(1-i\tau_{eq}\omega)^{2}\mathcal{S}^{(a)}(x;\omega,\vec{k})}{2(i\tau_{eq}\,k)^{3}\chi(x)}+\frac{i\mathcal{T}^{(a)}(x;\omega,\vec{k})}{2k\tau_{eq}\chi(x)}\right).$ (27d)
where $a=(s,l)$. For a massless system of particles, i.e. $x\to 0$, we obtain the previous results [27].
@@ III.1 Hydrodynamic limit
$G^{(s)}_{00}(x;\omega,\vec{k})=\chi(x)\frac{\mathcal{N}^{0(s)}(x;\omega,\vec{k})}{\mathcal{D}^{0(s)}(x;\omega,\vec{k})},$ (28)
@@
$\mathcal{N}^{0(s)}(x;\omega,\vec{k})=\frac{i(1-i)(1-i\tau_{eq}\omega)}{\tau_{eq}\omega}-\frac{ik^{2}\tau_{eq}N(x)}{15\omega(1-i\tau_{eq}\omega)K_{2}(x)\left(x^{2}K_{0}(x)+xK_{1}(x)-K_{2}(x)\right)},$
$\mathcal{D}^{0(s)}(x;\omega,\vec{k})=1-\frac{ik^{2}\tau_{eq}D(x)}{15\omega(1-i\tau_{eq}\omega)^{2}K_{2}(x)\left(x^{2}K_{0}(x)+xK_{1}(x)-K_{2}(x)\right)},$ (29)
@@
The lowest lying modes are solutions of spectral curves $\mathcal{D}^{(l)}(x;\omega,\vec{k})=0$ at small momenta which are obtained as follows
$\omega^{(nh)}_{(1,2)}\bigg{|}_{x\ll 1}$ $=-\frac{i}{\tau_{eq}}\pm\frac{2k}{\sqrt{15}}\left(1-\frac{x^{2}}{4}+\cdots\right)+\frac{ik^{2}\tau_{eq}}{6}\left(1-\frac{x^{2}}{2}+\cdots\right)+\mathcal{O}(k^{3}),$
$\omega^{(h)}\bigg{|}_{x\ll 1}$ $=-\frac{ik^{2}\tau_{eq}}{3}\left(1-\frac{x^{2}}{2}+\cdots\right)+\mathcal{O}(k^{4}),$ (31)
@@
$G^{(l)}_{00}(x;\omega,\vec{k})\bigg{|}_{x\gg 1}=\chi(x)\frac{\mathcal{N}^{0(l)}(x;\omega,\vec{k})}{\mathcal{D}^{0(l)}(x;\omega,\vec{k})},$ (32)
where
$\mathcal{N}^{0(l)}(x;\omega,\vec{k})=k^{2}\tau_{eq}^{2}(1-i\tau_{eq}\omega)\left(xK_{3}(x)-5K_{4}(x)\right)^{2},$ (33)
$\mathcal{D}^{0(l)}(x;\omega,\vec{k})=k^{2}\tau_{eq}^{2}\left(xK_{3}(x)-5K_{4}(x)\right)^{2}-i\tau_{eq}\omega x^{2}K_{2}(x)\bigg{(}xK_{3}(x)(1-i\tau_{eq}\omega)+(3k^{2}\tau_{eq}^{2}-5(1-i\tau_{eq}\omega)^{2})K_{4}(x)\bigg{)}.$
In the large $x$ limit, the lowest modes are given as follows
$\omega^{(nh)}_{(1,2)}\bigg{|}_{x\gg 1}$ $=-\frac{i}{\tau_{eq}}\pm k\sqrt{\frac{2}{x}}+\frac{ik^{2}\tau_{eq}}{2x}+\mathcal{O}(k^{3}),$
$\omega^{(h)}\bigg{|}_{x\gg 1}$ $=-\frac{ik^{2}\tau_{eq}}{x}+\mathcal{O}(k^{4}),$ (34)
This implies that for extremely massive quasiparticles, the hydro and non-hydro modes are stable. In Fig. 2 we plot the diffusive hydro modes for different cases. The left figure indicates the modes at $x=0.1$ for different $k\tau_{eq}$ values. The right figure shows the same results for $k\tau_{eq}=0.1$ for different $x$ values. It can be seen that the mass moves the poles into the unstable zone. This is because mass changes the diffusion constant, thereby influencing the location of hydro poles.
@@
$\lim_{\omega\rightarrow 0}\lim_{k\rightarrow 0}\frac{\omega}{k^{2}}{\rm Im}\,G_{00}(\omega,k)$ $=\chi D\,$
$\lim_{\omega\rightarrow 0}\lim_{k\rightarrow 0}\frac{1}{k^{2}}{\rm Re}\,G_{00}(\omega,k)$ $=\chi D\tau_{D}.$ (35)
For small $x$, using the expressions given in Eqs. (28) and (III.1) will readily obtain
$D(x)\bigg{|}_{x\ll 1}=\frac{\tau_{eq}}{3}\left(1-\frac{x^{2}}{2}+\cdots\right),$
$\tau_{D}\bigg{|}_{x\ll 1}=\tau_{eq}.$ (36)
For large $x$, the corresponding transport coefficients are
$D(x)\bigg{|}_{x\gg 1}=\tau_{eq}\left(\frac{1}{x}+\mathcal{O}(\frac{1}{x^{2}})\right),$
$\tau_{D}\bigg{|}_{x\gg 1}=\tau_{eq}.$ (37)
@@ IV Energy-Momentum Correlators
$\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)=\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}=\frac{T_{0}^{4}x^{3}}{2\pi^{2}}\bigg{(}xK_{2}(x)+3K_{3}(x)\bigg{)},$
@@
$\delta T^{0i}=-\frac{\tau_{eq}}{d}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{0i}(\omega,\vec{k})\frac{\mathcal{I}_{4}(v,\omega,k)}{i\tau_{eq}\omega}+\Gamma^{0}_{3i}(\omega,\vec{k})\mathcal{I}_{6}(v,\omega,k)\bigg{\}},\,\,i=1,2,$
$\delta T^{33}=\frac{\delta T^{00}c_{s}^{2}(1-i\tau_{eq}\omega)}{k^{2}\tau_{eq}^{2}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,k)+\frac{\delta T^{03}}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{5}(v,\omega,k)$
$\hskip 21.33955pt-\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{00}(\omega,\vec{k})\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}}\mathcal{I}_{1}(v,\omega,k)+2\,\Gamma^{0}_{03}(\omega,\vec{k})\mathcal{I}_{5}(v,\omega,k)+\Gamma^{0}_{33}(\omega,\vec{k})\mathcal{I}_{7}(v,\omega,k)$
$\hskip 142.26378pt+(\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\mathcal{I}_{8}(v,\omega,k)\bigg{\}},$
$\delta T^{ii}=\frac{\delta T^{00}c_{s}^{2}}{i\tau_{eq}\omega(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{4}(v,\omega,k)+\frac{\delta T^{03}}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,k)$
$\hskip 21.33955pt-\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{00}(\omega,\vec{k})\frac{\mathcal{I}_{4}(v,\omega,k)}{i\tau_{eq}\omega}+2\,\Gamma^{0}_{03}(\omega,\vec{k})\mathcal{I}_{6}(v,\omega,k)+\Gamma^{0}_{33}(\omega,\vec{k})\mathcal{I}_{8}(v,\omega,k)$
$\hskip 142.26378pt+(2\Gamma^{0}_{ii}(\omega,\vec{k})+\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\mathcal{I}_{9}(v,\omega,k)\bigg{\}},\,\,\,i=1,2,$
$\delta T^{12}=-\,\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\Gamma^{0}_{12}(\omega,\vec{k})\mathcal{I}_{9}(v,\omega,k),$
$\delta T^{3i}=\frac{\delta T^{0i}}{\varepsilon_{eq}+P_{eq}}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,k)$
$\hskip 21.33955pt-\,\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{0i}(\omega,\vec{k})\mathcal{I}_{6}(v,\omega,k)+\Gamma^{0}_{3i}(\omega,\vec{k})\mathcal{I}_{8}(v,\omega,k)\bigg{\}},\,\,i=1,2.$
@@
$G^{00,00}(\omega,k)=-\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\bigg{(}1+\frac{i\,b\,c_{s}^{2}\,k\,\tau_{eq}}{ac-b^{2}c_{s}^{2}}\bigg{)},$
$G^{00,03}(\omega,k)=\frac{\omega}{k}\left(G^{00,00}(\omega,k)+\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\right),$
$G^{00,33}(\omega,k)=\frac{\omega}{k}G^{00,03}(\omega,k),$
$G^{00,11}(\omega,k)=G^{00,22}(\omega,k)$
@@
$G^{03,00}(\omega,k)=\frac{\omega}{k}\left(G^{00,00}(\omega,k)+\frac{\varepsilon_{eq}+P_{eq}}{c_{s}^{2}}\right)=-(\varepsilon_{eq}+P_{eq})\,\frac{\,i\,\tau_{eq}\,\omega\,b}{ac-b^{2}c_{s}^{2}}=G^{00,03}(\omega,k),$
$G^{03,03}(\omega,k)=\frac{\omega}{k}G^{00,03}(\omega,k)-\left(\varepsilon_{eq}+P_{eq}\right),$
$G^{03,33}(\omega,k)=\frac{\omega}{k}G^{03,03}(\omega,k),$
$G^{03,11}(\omega,k)=G^{03,22}(\omega,k)=\frac{\omega}{k}\left(G^{00,11}(\omega,k)-\left(\varepsilon_{eq}+P_{eq}\right)\right),$
$G^{0i,0i}(\omega,k)=\frac{k}{\omega}G^{0i,3i}(\omega,k)=\frac{-i\tau_{eq}\,k}{d}\,\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{6}(v,\omega,k),\qquad i=1,2,$
$G^{12,12}(\omega,k)=-i\tau_{eq}\omega\,\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{9}(v,\omega,k),$
$G^{3i,0i}(\omega,k)=\frac{k}{\omega}G^{3i,3i}(\omega,k)=\frac{\omega}{k}G^{0i,0i}(\omega,k).$ (48)
It is evident that the Ward identity relates different two-point functions and the only independent components are $(G^{00,00},G^{00,11},G^{01,01},G^{12,12})$. To proceed, we give a detailed expansion of the $b$ parameter in the small and large $x$ limits
$b\bigg{|}_{x\ll 1}$ $=\frac{-i}{k\tau_{eq}c_{s}^{2}}\bigg{\{}1+\frac{i\omega\tau_{eq}-1}{2ik\tau_{eq}}\bigg{(}L(\omega,k)+g_{1}(x)\left(L(\omega,k)-\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$\hskip 49.79231pt-g_{2}(x)\left(L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(k^{2}\tau_{eq}^{2}-(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\mathcal{O}(\frac{x^{6}}{\tilde{p}^{6}})\bigg{)}\bigg{\}},$
$b\bigg{|}_{x\gg 1}$ $=-\frac{ik\tau_{eq}}{(1-i\tau_{eq}\omega)^{2}}\bigg{(}h_{1}(x)-h_{2}(x)\frac{k\tau_{eq}\left(3k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)}{(1-i\tau_{eq}\omega)^{2}}\bigg{)},$
@@
$g_{1}(x)\equiv\frac{x\left(xK_{1}(x)+K_{2}(x)\right)}{2\left(xK_{2}(x)+3K_{3}(x)\right)},\qquad g_{2}(x)\equiv\frac{x^{3}\left(K_{2}(x)-3K_{0}(x)\right)}{16\left(xK_{2}(x)+3K_{3}(x)\right)},$
$h_{1}(x)\equiv\frac{5xK_{2}(x)+(30+x^{2})K_{3}(x)}{x^{2}K_{3}(x)},\qquad h_{2}(x)\equiv\frac{7xK_{3}(x)+(56+x^{2})K_{4}(x)}{x^{3}K_{3}(x)}.$ (50)